OpenStudy (anonymous):

Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. e^2x-9e^x-20=2

OpenStudy (anonymous):

start with \[e^{2x}-9e^x-22=0\] then solve the quadratic equation

OpenStudy (anonymous):

\[(e^x-11)(e^x+2)=0\] \[e^x-11=0\] \[e^x=11\] \[x=\ln(11)\] \[e^x+2=0\] \[e^x=-2\] impossible since \[e^x>0\]